The measure of one interior angle of a parallelogram is 3.5 times the measure of another angle. What is the measure of each angle?
Answer by ikleyn(52817)

.

x + 3.5x = 180  degrees


    4.5x = 180

       x  = 180/4.5 = 40.


The angles are  40 degrees and  140 degrees.

solved.